Two long parallel wires are 10 cm apart and carry currents of 6 A and 4 A. what is the field B at 4 A carrying wire due to 6 A wire.
1
Expert's answer
2020-05-12T09:56:07-0400

"B=\frac{\mu_oI}{2\pi R}=\frac{4\pi\times10^{-7}\times6}{2\pi\times0.1}=12\times10^{-6}H"